OpenStudy (anonymous):

Solve the problem. When making a long distance call from a certain pay phone, the first three minutes of a call cost $1.10. After that, each additional minute or portion of a minute of that call costs $0.30. Use an inequality to find the number of minutes one can call long distance for $4.70. A) 15 minutes or fewer B) 4 minutes or fewer C) 12 minutes or fewer D) 16 minutes or fewer

OpenStudy (anonymous):

3(x+.30)=4.70?

hero (hero):

It says to use an inequality

OpenStudy (anonymous):

eww! lol k

hero (hero):

From where did you get that formula? Did you make that up?

OpenStudy (anonymous):

yeah i like trying to put them together in differn orders not even close huh?

hero (hero):

No because you need to basically use graph paper to do this. You need to plot the points (3, 1.10) and (4, 1.40) then use y = mx + b again.

hero (hero):

I'm not even going to do anything else. You need to figure this out yourself.

OpenStudy (anonymous):

so is the slope-30/1

OpenStudy (anonymous):

um no its to hard

hero (hero):

Hardly: y = mx + b 1.10 = 3m + b 1.40 = 4m + b b = 1.10 - 3m b = 1.40 - 4m b = b 1.10 - 3m = 1.40 - 4m 4m - 3m = 1.40 - 1.10 m = .30 b = 1.10 - 3m b = 1.10 - 3(.30) b = 1.10 - .90 b = .20 y = .30x + .20

hero (hero):

Now we need to find the number of minutes you can call for 4.70. To do that you have to make an inequality

hero (hero):

4.70 ≥ .30x + .20 4.70 - .20 ≥ .30x 4.50 ≥ .30x 4.50/.30 ≥ x 15 ≥ x

hero (hero):

So 15 min or fewer

hero (hero):

You seem to be in disbelief regarding the simplicity of this problem.

OpenStudy (anonymous):

no im just extremely exhuasted i know these its just i cant remember Im takinga pretest so i can get in to my midterms at the school. I have 76 math problems. and I need a 70 to pass I took it once and failed so now im here trying to pass it so i can take the darn test

hero (hero):

You cannot guess these things. You have to learn the concepts or else you will fail everytime.

OpenStudy (anonymous):

i need to know if they are right

OpenStudy (anonymous):

i do its just i have a science test to study for to and my children are sick im sick i just want to get it over with

hero (hero):

It's not like I have plenty of free time on my hands either. I'm actually taking a break from the work I have to do. I can't do 76 problems.

OpenStudy (anonymous):

i know im not asking and i really do appreciate your help